Find three consecutive integers such that the sum of twice the smallest and three times the largest is 126. Please explain how to do it thanks:)
Answer by stanbon(75887)

Find three consecutive integers such that the sum of twice the smallest and three times the largest is 126. Please explain how to do it
---------------
1st: x
2nd: x+1
3rd: x+2
----
Equation:
2x+3(x+2) = 126
---
2x+3x+6 = 126
5x = 120
x = 24
x+1 = 25
x+2 = 26
====================
Cheers,
Stan H.